# Simulation of time crystal behavior for a few boson chiral soliton model in a ring

### v3

@@ III.3 Time crystal dynamics from measurement
The basic message from our two-particle simulations is that using a weak measurement of particle position can lead to time crystal behavior for the remaining particle on the ring. This is the case since the weak measurement can dominantly excite a non-dispersive and rotating solution on the ring, whereas the more precise position measurement excites a broader wavepacket of the ring modes that leads to the observed decay and lack of time crystal behavior. The benefit of the two-particle case is that it reveals the role played by the position measurement uncertainty in a transparent manner.
@@ IV.1 Three-particle Schrödinger equation
$i{\partial\Psi\over\partial t}=\sum_{j=1}^{3}\left(-i{\partial\over\partial\theta_{j}}-{\ell\over 2}-\kappa\rho(\theta_{1},\theta_{2},\theta_{3}))\right)^{2}\Psi+{g\over 2}\rho(\theta_{1},\theta_{2},\theta_{3})\Psi.$ (17)
@@
$i{\partial\Psi\over\partial t}=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}+{1\over 3}{\partial^{2}\over\partial s^{2}}\right)\Psi+2i\kappa\rho(x,y){\partial\Psi\over\partial s}+3\left[{\ell\over 2}+\kappa\rho(x,y)\right]^{2}\Psi+{g\over 2}\rho(x,y)\Psi.$ (22)
@@ IV.2 Three-particle ground state
$\varepsilon\varphi(x,y)=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}\right)\varphi+3\left[{p\over 3}-{\ell\over 2}-\kappa\rho(x,y)\right]^{2}\varphi+{g\over 2}\rho(x,y)\varphi,$ (24)
@@
We have solved Eq. (24) numerically for a range of parameters using the imaginary time method LehToiElo2007 , in particular to obtain the ground state $\varphi(x,y)$ and associated scaled energy $\varepsilon$. Figure 5 shows numerical results for parameters $\ell=2,g=-0.8,\kappa=0.1,q=50$. Specifically, Fig. 5(a) shows the scaled energy versus $p$ and reveals that the ground state occurs when ${p\over N}={\ell\over 2}$, or $p=3$ for this example, and (b) shows a line plot of the ground state profile $|\varphi(x,y)|^{2}$ versus $x$ and $y$ for $p=3$.
Figure 5: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=3$. Parameters are $\ell=2,g=-0.8,\kappa=0.1,q=50$.
IV.3 Time crystal dynamics from measurement
To proceed we first recognize that the ground state represented by the probability density in Fig. 5(b) does not correspond to a localized state on the ring since $|\varphi(x,y)|^{2}$ depends only on the Jacobi coordinates $x$ and $y$, and there is no information concerning the COM position $s$ of the particle. We can also reconstruct the ground state three-body wave function for one coordinate fixed, say $\theta_{3}=0$, and $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$ (at $t=0$) is shown in Fig. 6(a): We note that the width of this ground state density profile is around $w\sim 1$, that provides a characteristic scale for localized three-particle states that may appear in the system.
@@ IV.3 Time crystal dynamics from measurement
We next extend the measurement model to the case with three initial particles. In particular, we assume that the position of particle $3$ is measured yielding a value $\theta_{3}^{(0)}$, but allow for the measurement to be imprecise, so that the wave function for the remaining particles may be written as
@@
where the function $G(\theta_{3}-\theta_{3}^{(0)})$ reflects the uncertainty of the position measurement, and we again use the model in Eq. (12). After the measurement the wave function $\psi(\theta_{1},\theta_{2},t=0)$ (suitably normalized) for the remaining particles is given by Eq. (26), and subsequent time development is governed by the two-particle Schrödinger equation
$i{\partial\psi\over\partial t}=-\left({\partial^{2}\over\partial\theta_{1}^{2}}+{\partial^{2}\over\partial\theta_{2}^{2}}\right)\psi+2i\left({\ell\over 2}+\kappa\rho\right)\left({\partial\over\partial\theta_{1}}+{\partial\over\partial\theta_{2}}\right)\psi+2\left({\ell\over 2}+\kappa\rho\right)^{2}\psi+{g\over 2}\rho\psi,$ (27)
@@
Figure 6(b) shows a simulation of the quantum dynamics according to Eq. (27) following a precise measurement for which we set $\theta_{3}^{(0)}=0$ and $n=50$ so that $\Delta\theta\sim 0.2$. In this case the uncertainty is much less than the expected width $w\sim 1$ for a three-particle localized state, so the measurement is expected to be very disruptive. The figure shows a color coded plot of the probability density for measuring a second particle
@@
versus $t$ and $\theta_{1}$. Then, for example, if we select $\theta_{1}=\theta_{3}^{(0)}=0$, and track the probability density in time along the horizontal axis, we see what might be a periodic variation indicative of a time crystal. But this is an illusion based on us choosing $\theta_{1}=\theta_{3}^{(0)}$, a quantity we cannot know a priori, and if we chose another $\theta_{1}$ at which to monitor the time variation of the probability density we would generally find another variation with time. Another way to see this from Fig. 6(b) is that the spatial density profile $|\psi(\theta_{1},t=0)|^{2}$ becomes rapidly distorted as time increases. This is in perfect keeping with the findings of SKS that too precise a position measurement of one particle can produce an localized state initially but this will become rapidly dispersed before the initial localized solution can exhibit a single rotation around the ring.
@@
Figure 6: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$, and (b) simulation for a measurement with $n=50$ that localizes particle $3$ with an uncertainty $\Delta\theta\sim 0.2$. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it does not reflect time crystal behavior. Parameters are the same as as in Fig. 5.
@@
For time crystal behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. We may use our previous estimate for the magnitude of the scaled rotation rate $|v|\leq{2|p|\over N}$, or $|v|\leq 2$ for the current example. Figure 7 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 6(b): In this case the uncertainty is greater than the expected width $w\sim 1$ for a three-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates with velocity $v\sim-1$ Neq3 , and it persists over several rotations around the ring. In addition, if one selects a value for $\theta_{1}$ in Fig. 7 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior to a good approximation.
@@
To summarise, we deduce that the initial measurement of particle $3$ leads to spontaneous formation of a localized state that is a two-particle analogue of a mean-field chiral soliton Jackiw , and that this two-particle soliton can continue rotating whilst intact over several rotation periods. This can happen here since we considered a weak measurement as opposed to the precise measurement that leads to decay of the rotating soliton, see Fig. 6(b).
@@
Figure 7: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining two particles after the measurement of particle $3$ to the two-particle limit of the chiral soliton. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as as in Fig. 5.
Figure 8: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=6$. Parameters are $\ell=4,g=-1,\kappa=0.1,q=50$.
@@ V Summary and conclusions
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 , and this proposal met with some criticism in the literature SKS_EPL ,SKS_comment -OW_lack . The goal of the present paper was to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time crystal behavior is possible in the context of the type of calculation in Ref. SKS_EPL . In particular, we initialize the system in the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson localized state (soliton for $N=3$) can persist over several revolutions of the ring as required for time crystal behavior. We find that for an imprecise or weak position measurements time crystal behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS_EPL .
@@
One feature from the simulations is that in the presence of chirality the quantum ground state of the system can carry orbital angular momentum pPlanck-constant-over-2-$\hbar p$, with $p$ the center-of-mass winding number which can be non-zero for the state of lowest energy. This is illustrated in Fig. 2(a) for $N=2$ with $p=2$ for minimum energy, and Fig. 5(a) for $N=3$ with $p=3$, both for $\ell=2$. More generally, we find that for minimum energy ${p\over N}={\ell\over 2}$, and this is illustrated in Fig. 8(a) for $\ell=4$, giving $p=6$ for minimum energy. The point is that the presence of chirality and non-zero orbital angular momentum appear to be key requirements for time-crystal behavior to be a possibility in our model, the imprecise measurement facilitating transfer of the orbital angular momentum to the center-of-mass momentum of the localized soliton. Estimates for the rotation rates were given but details of this process remain to be elucidated.
@@
Clearly the small particle numbers and parameters used here are not compatible with experiments, but are rather intended to show that time crystal behavior is not ruled out in our model. We remark that the same behavior could be seen for other combinations of parameters as long as a weak measurement is employed, with $\Delta\theta>w$. Restricting to a few bosons allowed for numerical simulations to be performed for our chiral soliton model without undue approximations, but our conclusions may still be of relevance in the more general case with large $N$. According to the Wilczek model analyzed in Ref. SKS_EPL the soliton lifetime $t_{c}$ is expected to scale with $N$. If this scaling also applies to our chiral model, then if $N=3$ can show time crystal behavior it should be expected for larger $N$ as the lifetime increases.
Finally, there is the issue of whether our simulations point to genuine time crystal behavior. This would require that the periodic motion seen in Fig. 7 would persist for all times for $N=3$, and we cannot be assured that this is the case based on numerical solutions. What we can say for sure is the time crystal behavior can persist over several revolutions of the ring if imprecise measurements are allowed for. In future publications we plan to extend this initial numerical study to larger particle numbers, both numerically and analytically, and thereby move the project towards experimentally accessible parameter ranges.

### v4

@@ III.4 Role of the measurement strength
The basic message from our two-particle simulations is that using a weak measurement of particle position time-crystal-like behavior is possible for the remaining particle on the ring if $\Delta\theta/w>1$. This is the case since the weak measurement can dominantly excite a non-dispersive and rotating solution on the ring, whereas the more precise position measurement excites a broader wavepacket of the ring modes that leads to the observed loss of contrast and lack of time-crystal-like behavior. The benefit of the two-particle case is that it reveals the role played by the position measurement uncertainty in a transparent manner.
@@ IV.1 Three-particle Schrödinger equation
$i{\partial\Psi\over\partial t}$ $=$ $\sum_{j=1}^{3}\left(-i{\partial\over\partial\theta_{j}}-{\ell\over 2}-\kappa\rho(\theta_{1},\theta_{2},\theta_{3}))\right)^{2}\Psi$ (19)
$+{g\over 2}\rho(\theta_{1},\theta_{2},\theta_{3})\Psi.$
@@
$i{\partial\Psi\over\partial t}$ $=$ $-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}+{1\over 3}{\partial^{2}\over\partial s^{2}}\right)\Psi+2i\kappa\rho(x,y){\partial\Psi\over\partial s}$ (24)
$+3\left[{\ell\over 2}+\kappa\rho(x,y)\right]^{2}\Psi+{g\over 2}\rho(x,y)\Psi.$
@@ IV.2 Three-particle ground state
$\varepsilon\varphi(x,y)=-\bigg{(}2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}\bigg{)}\varphi$
$+3\bigg{(}{p\over 3}-{\ell\over 2}-\kappa\rho(x,y)\bigg{)}^{2}\varphi+{g\over 2}\rho(x,y)\varphi,$ (26)
@@
We have solved Eq. (IV.2) numerically for a range of parameters using the imaginary time method LehToiElo2007 , in particular to obtain the ground state $\varphi(x,y)$ and associated scaled energy $\varepsilon$. Figure 6 shows numerical results for parameters $\ell=2,g=-0.7,\kappa=0.05,q=50$. Specifically, Fig. 6(a) shows the scaled energy versus $p$ and reveals that the ground state occurs when ${p\over N}={\ell\over 2}$, or $p=3$ for this example, and (b) shows a line plot of the ground state profile $|\varphi(x,y)|^{2}$ versus $x$ and $y$ for $p=3$.
Figure 6: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=3$. Parameters are $\ell=2,g=-0.7,\kappa=0.05,q=50$.
IV.3 Time-crystal-like dynamics from measurement
To proceed we first recognize that the ground state represented by the probability density in Fig. 6(b) does not correspond to a localized state on the ring since $|\varphi(x,y)|^{2}$ depends only on the Jacobi coordinates $x$ and $y$, and there is no information concerning the COM position $s$ of the particle. Again, since the ground state has a winding number $p\neq 0$ it does represents a persistent flow RyuAndCla07 . We can also reconstruct the ground state three-body wave function for one coordinate fixed, say $\theta_{3}=0$, and $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$ (at $t=0$) is shown in Fig. 7(a). We note that the width (full-width at half-maximum) of this ground state density profile is $w\sim 1.4$, that provides a characteristic scale for localized three-particle states that may appear in the system.
@@ IV.3 Time-crystal-like dynamics from measurement
We next extend the measurement model to the case with $N=3$ particles. In particular, we assume that the position of particle $3$ is measured yielding a value $\theta_{3}^{(0)}$, but allow for the measurement to be imprecise, so that the wave function for the remaining particles may be written as
$\psi(\theta_{1},\theta_{2},t=0)=\Psi(\theta_{1},\theta_{2},\theta_{3}^{(0)})=$
@@
where the function $G(\theta_{3}-\theta_{3}^{(0)})$ reflects the uncertainty of the position measurement, and we again use the model in Eq. (12). After the measurement the wave function $\psi(\theta_{1},\theta_{2},t=0)$ (suitably normalized) for the remaining particles is given by Eq. (IV.3), and subsequent time development is governed by the two-particle Schrödinger equation
$i{\partial\psi\over\partial t}$ $=$ $-\bigg{(}{\partial^{2}\over\partial\theta_{1}^{2}}+{\partial^{2}\over\partial\theta_{2}^{2}}\bigg{)}\psi$ (29)
$+2i\bigg{(}{\ell\over 2}+\kappa\rho\bigg{)}\bigg{(}{\partial\over\partial\theta_{1}}+{\partial\over\partial\theta_{2}}\bigg{)}\psi$
$+2\bigg{(}{\ell\over 2}+\kappa\rho\bigg{)}^{2}\psi+{g\over 2}\rho\psi,$
@@
Figure 7(b) shows a simulation of the quantum dynamics according to Eq. (29) following a precise measurement for which we set $\theta_{3}^{(0)}=0$ without loss of generality, and $n=50$ so that $\Delta\theta\sim 0.34$ (full-width at half-maximum of $|G(\theta)|^{2}$): As for the $N=2$ case, we chose $n=50$ as this approaches a precise measurement as closely as possible within our non-local model with $q=50$ in Eq. (12). In particular, Fig. 7(b) shows a color-coded plot of the single-particle probability density for measuring a second particle
@@
In this case the uncertainty is much less than the expected width $w\sim 1.4$ for a three-particle localized state, so the measurement is expected to be very disruptive. We note that although the probability density appears to be varying periodic in time with period $2\pi$, for $t>4\pi$ the profile is seen to deviate significantly from this periodicity. Moreover, as expected for a strong measurement, the initial localized solution evident in the single-particle probability density $p(\theta_{1},t=0)$ becomes rapidly distorted as time increases, not simply displaced as expected for a time crystal, so this example does not display time-crystal-like behavior. This is in perfect keeping with the findings of SKS that too precise a position measurement of one particle can produce a localized state initially but this will become rapidly dispersed before the initial localized solution can exhibit even a single rotation around the ring.
@@
Figure 7: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$, and (b) simulation of the single-particle probability density $p(\theta_{1},t)$ for a measurement with $n=50$ that localizes particle $3$ with an uncertainty $\Delta\theta\sim 0.34$. This example has $\Delta\theta/w=0.24$ and does not display time-crystal-like behavior. Parameters are the same as as in Fig. 6.
@@
For time-crystal-like behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. We may use our previous estimate for the magnitude of the scaled rotation rate $|v|\leq{2|p|\over N}$, or $|v|\leq 2$ for the current example. Figure 8 shows an example of time-crystal-like behavior using $n=1$ and $\Delta\theta\sim 2.3$, which means a much weaker measurement than in Figure 7(b). In this case the uncertainty is greater than the expected width $w\sim 1.4$ for a three-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates with velocity $v\sim-1$ Neq3 , and in the figure is seen to persist over $>10$ rotations around the ring. This example therefore exhibits time-crystal-like behavior to a good approximation.
@@
To summarise, we deduce that the initial measurement of particle $3$ leads to spontaneous formation of a localized state that is a two-particle analogue of a mean-field chiral soliton Jackiw , and that this two-particle soliton can continue rotating largely intact over several rotation periods. This can happen here since we considered a weak measurement as opposed to the precise measurement that leads to decay of the rotating soliton, see Fig. 7(b).
@@
Figure 8: Simulation of the single-particle probability density $p(\theta_{1},t)$ for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2.3$ that couples the remaining two particles after the measurement of particle $3$ to the two-particle limit of the chiral soliton. This example has $\Delta\theta/w=1.63$ and does display time-crystal-like behavior. Parameters are the same as as in Fig. 6.
IV.4 Role of the measurement strength
The results shown in Figs. (7) and (8) show the extremes of strong and weak measurements, respectively, and how they impact the quantum dynamics for $N=3$ particles as reflected in the probability density $p(\theta_{1},t)$ of the system. We next provide some details of how the strength of the measurement impacts whether time-crystal-like behavior is present or not.
Table 1: Data on measurement strength
n $\Delta\theta/w$ $t_{c}/2\pi$
1 1.63 $>50$
2 1.16 1.7
3 0.96 0.7
4 0.83 0.2
5 0.74 0.2
6 0.68 0.2
Turning first to the case with a strong measurement in Fig. 7(b) with $n=50$ and $\Delta\theta/w=0.24$, we note that the single-particle probability density $p(\theta_{1},t)$ is no longer periodic in time. This is the case since the quantum dynamics after the measurement is now governed by the two-dimensional Schrödinger equation (29) which will have a more complex eigenmode spectrum than the linear Eq. (13) for the case of $N=2$ particles. As a result the wave packet dynamics for the case of $N=3$ particles can be far richer. In particular, the contrast $C(t)$ defined in Eq. (17) can now get progressively smaller as time increases, though recurrences may still occur in general. Indeed, SKS SKS_EPL defined the lifetime $t_{c}$ of the revolving solution as the first time for which the normalized contrast $C(t_{c})/C(0)=0.5$ drops to one-half its initial value. We adopt this approach here, and for the example in Fig. 7(b) $t_{c}\simeq 0.2$ meaning that the initial localized solution executes only a fraction of a revolution of the ring that takes time $2\pi$. In contrast, in Fig. 8 with $n=1$ and $\Delta\theta/w=1.63$, time-crystal-like behavior is clear over the duration $t=[0,80]$ of the simulation, which is greater than $t_{c}/2\pi>10$ revolutions of the ring, and the normalized contrast remains $>0.9$ throughout. We limited the maximum time to $80$ only so the features in Fig. 8 were resolvable. Indeed, for times over which we can maintain numerical accuracy, which amounts to more than $t_{c}/2\pi>50$ revolutions, the normalized contrast remains well above $0.5$. Within numerical accuracy it is therefore legitimate to claim that this $N=3$ example clearly shows time-crystal-like behavior, though we cannot claim that it persists indefinitely.
To explore how the time-crystal-like behavior varies with the strength of the measurement we performed simulations for a range of different values of $n$ in Eq, (12), and some results are tabulated in Table 1. The left column gives $n$, the center column gives the corresponding value for $\Delta\theta/w$, and the right hand column gives $t_{c}/2\pi$ which provides an estimate of the number of revolutions executed by the solution before it decays. We see that for $n\geq 4$ the solution executes only a fraction of a revolution, which is consistent with the fact that these correspond to strong measurements with $\Delta\theta/w<1$. In contrast, for $n=2,3$ for which $\Delta\theta/w\simeq 1$, the solution executes roughly $1-2$ revolutions before decaying. Finally, for $n=1$ and $\Delta\theta/w=1.63$ we have the example shown in Fig. 8 where the lifetime increases dramatically to $>50$, indicative of a divergence. This apparent divergence in the lifetime signals that the time-crystal-like behavior can persist for a large number of revolutions of the ring.
@@ V Summary and conclusions
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 , and this proposal met with some criticism in the literature SKS_EPL ,SKS_comment -OW_lack . The goal of the present paper was to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time-crystal-like behavior is possible in the context of the type of calculation in Ref. SKS_EPL . In particular, we initialize the system in the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson localized state (soliton for $N=3$) can persist over several revolutions of the ring as required for time-crystal-like behavior. We find that for an imprecise or weak position measurements time-crystal-like behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS_EPL .
@@
One feature from the simulations is that in the presence of chirality the quantum ground state of the system can carry orbital angular momentum pPlanck-constant-over-2-$\hbar p$, with $p$ the center-of-mass winding number. This means that the lowest energy state can be a persistent flow RyuAndCla07 . The point is that the presence of chirality and a ground state that is a persistent flow appear to be key requirements for time-crystal behavior to be a possibility in our model, the imprecise measurement facilitating transfer of the orbital angular momentum to the center-of-mass momentum of the localized soliton. Estimates for the rotation rates were given but details of this process remain to be elucidated.
@@
Clearly the small particle numbers and parameters used here are not compatible with experiments, but are rather intended to show that time-crystal-like behavior is not ruled out in our model. We remark that the same behavior could be seen for other combinations of parameters as long as a weak measurement is employed, with $\Delta\theta/w>1$. Restricting to a few bosons allowed for numerical simulations to be performed for our chiral soliton model without undue approximations, but our conclusions may still be of relevance in the more general case with large $N$. According to the Wilczek model analyzed in Ref. SKS_EPL the soliton lifetime $t_{c}$ is expected to scale with $N$. If this scaling also applies to our chiral model, then if $N=3$ can show time-crystal-like behavior it should be expected for larger $N$ as the lifetime increases. To explore the scaling of the lifetime with $N$ we calculated the root-mean-square width $W(t)$ of the evolving solution after a weak measurement $(n=1)$ over the time range $t=[0,2\pi]$, and Fig. 9 shows $W(t)/W(0)$ for $N=2$ and $N=3$ for parameters $\kappa=0.2,g=-2.0$ (similar results were found for the parameters $\kappa=0.05,g=-0.7$). In particular, the plot shows that the width $W(t)$ initially grows more slowly for the case $N=3$ compared to $N=2$, and this adds some credibility to the statement that “the soliton lifetime $t_{c}$ is expected to scale with $N$.” The reversal of the initial increase in $W(t)$ is a consequence of the recurrence that occurs due to the periodic boundary conditions imposed by the ring, this leading to the fact that for $N=2$ the solution must be strictly periodic in time with period $2\pi$. In future publications we plan to extend this initial numerical study to larger particle numbers, both numerically and analytically, and thereby move the project towards experimentally accessible parameter ranges.
Figure 9: The plot shows the normalized root-mean-square width $W(t)/W(0)$ after a weak measurement $(n=1)$ versus $t$ for $N=2$ and $N=3$ for parameters $\kappa=0.2,g=-2.0$.